# Improving Detection of Watermarked Language Models

### v1

@@ 1 Introduction
The aim of watermarking is to bias the model so that first-party detection becomes more tractable. Most schemes do not modify the LLM’s training procedure but instead inject the watermark signal at inference time as part of the autoregressive decoding loop (Aaronson, 2023; Bahri and Wieting, 2024; Kuditipudi et al., 2023; Kirchenbauer et al., 2023; Dathathri et al., 2024). Meanwhile, non-watermark detection has mostly been viewed as a binary classification task, with the goal of discriminating one class (usually human-written text) from another (usually one or more models). Strategies here include training a binary classifier on labeled text samples or computing uncertainty-based scores, such as likelihood, under a specific LLM (Zellers et al., 2019; Solaiman et al., 2019; Gehrmann et al., 2019; Su et al., 2023; Mitchell et al., 2023). The focus of our work is on improving first-party detection by intelligently combining watermark- and non-watermark-based detection approaches.
The entropy of an LLM’s distribution over possible responses conditioned on a specific prompt is a key quantity in watermark detection. For example, Bahri and Wieting (2024) provides a lower-bound on the detection ROC-AUC as a function of the entropy of the sampled next-token distribution. In all schemes, watermark performance improves with more available entropy. As a concrete example, consider two input prompts to a watermarked LLM: (1) “Where is Mount Whitney?” and (2) “Write me a dreamy haiku about the John Muir Trail in California.” The first prompt is a fact-based inquiry whereas the second is more open-ended in nature, so the entropy will likely be higher for the second prompt than for the first and the generated haiku will be easier to detect via the watermark. The role of entropy for non-watermark detection is less studied. We show how non-watermark detectors, when configured in a hybrid setup, can substantially bolster watermarking in low entropy regimes. We study how these two approaches to detection can be combined effectively for better predictive performance than either part as well as computational advantages, and we recommend practical algorithms for deployment.
@@ 2.1 Watermarking
Although watermarking, sometimes referred to as linguistic steganography, has a storied past, interest in its application to modern large language models grew substantially after the seminal works of Kirchenbauer et al. (2023) and Aaronson (2023). Many successful techniques employ pseudo-random functions (PRFs) and cryptographic hashes and are applied token-by-token during the autoregressive decoding loop. Kirchenbauer et al. (2023) uses a PRF to bias the next-token probability distribution so that certain tokens known only to the watermarker are made more probable. The degree of bias is a hyper-parameter that trades off text distortion and watermark strength. Aaronson (2023) proposes a clever strategy that selects the token in the next-token distribution that has a high PRF value and a high likelihood in a way that makes the process appear distortion-free. Kuditipudi et al. (2023) applies a scheme similar to Aaronson (2023) but to increase robustness to attacks such as paraphrasing, the pseudo-random numbers are determined by cycling through a secret, predetermined sequence of values rather than by -grams. Meanwhile, SynthID (Dathathri et al., 2024) samples a large number of tokens from the next-token distribution and pits them head-to-head in a series of tournaments, the winner of which is selected as the next token. Lee et al. (2023) adapts Kirchenbauer et al. (2023)’s scheme for the code-generation task by applying the watermark only at decoding steps that have sufficient entropy.
Black-box schemes that only require a way to sample sequences from the LLM have been devised (Yang et al., 2023a; Giboulot and Furon, 2024; Chang et al., 2024; Bahri and Wieting, 2024). Most notably, Bahri and Wieting (2024) uses elements from Aaronson (2023) to construct a general framework for black-box distortion-free watermarking, which can be used effectively when white-box access is available.
Ways of evading, spoofing, and even stealing watermarks have been extensively studied (Krishna et al., 2024; Zhang et al., 2023; Gu et al., 2023; Jovanović et al., 2024). To address the weakness of many schemes to attacks such as token substitution or paraphrasing, watermarking based on semantics or invariant features has also been proposed (Liu et al., 2023; Hou et al., 2023; Ren et al., 2023; Yoo et al., 2023). Fernandez et al. (2023) tests various watermarking schemes on classical NLP benchmarks and introduces new statistical tests — for example, they suggest skipping duplicate -grams during testing. Liang et al. (2024) conducts a comprehensive assessment of watermarking strategies, finding that incorporating a non-watermark RoBERTa-based detector like the way we do can help boost robustness to attacks. The purpose of our work is to generalize this focused insight by proposing and then carefully evaluating various hybrid schemes.
@@ 2.2 AGC Detection
A powerful paradigm to detect AGC is to finetune a pretrained LM on sizable datasets for classification (Zellers et al., 2019; Solaiman et al., 2019); notably, Hu et al. (2023) and Tian et al. (2023) do so via adversarial and positive-unlabeled (PU) learning respectively, while Tay et al. (2020) applies it to the task of discriminating different LM generators from one another.
A second camp of approaches focuses on capturing statistical patterns of AGC using little to no training data — such as the intrinsic dimensionality of generated text (Tulchinskii et al., 2024), perplexity / rank / log-rank (Gehrmann et al., 2019; Su et al., 2023), perplexity curvature (Mitchell et al., 2023), and -gram patterns (Yang et al., 2023b).
Mireshghallah et al. (2024) shows, interestingly, that small LM generators make for better zero-shot detectors than larger ones.
Mao et al. (2024) leverages the observation that LLMs change more words when tasked to rewrite human text than to rewrite AGC. Zhang et al. (2024) shows that existing detectors are inadequate for detecting AI-revised human-written text, text that a human wrote with assistance from an LLM. Interestingly, Russell et al. (2025) found that humans who use LLMs for writing tasks can outperform many commercial and open source detectors even in the presence of paraphrasing and humanization.
As usage of large language models is still nascent, it remains to be seen precisely how they are used in the wild, be it with good or nefarious intentions. Bahri et al. (2021) conducts a large-scale study to this end by running a GPT-2 detector on half a billion web pages.
@@ 3.1 Models, Datasets, Hyperparameters, and Compute
We consider two distinct settings: when is Gemma-7B-instruct (Team et al., 2024b)111https://huggingface.co/google/gemma-7b-it and is Mistral-7B-instruct (Jiang et al., 2023)222https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1 as well as the reverse assignment.
We use two test datasets: databricks-dolly-15k333https://huggingface.co/datasets/databricks/databricks-dolly-15k (Conover et al., 2023), an open source dataset of instruction-following examples for brainstorming, classification, closed QA, generation, information extraction, open QA, and summarization. Exactly replicating Bahri and Wieting (2024), we use prompts from the brainstorming, generation, open QA (i.e. general QA), and summarization categories, whose human responses are at least 50 tokens long (save one example, which was removed because the prompt was extremely long); this amounts of 5,233 prompts total.
We use the training and test splits of eli5-category444https://huggingface.co/datasets/rexarski/eli5_category (Fan et al., 2019). Prompts are formed by concatenating the prompt and selftitle fields. Only examples whose prompt field is non-empty and contains a ? are kept — for a total of 83,089 train and 4,855 test samples.
@@ 3.3 Watermarks
Aaronson. At each step, Aaronson (2023) computes a pseudorandom number for each token as, , where is the preceding -gram, and denotes concatenation. Token is selected, where . At test time, -grams are extracted from the test query and the detection score is given by,
Bahri and Wieting (2024) notes that is not length-aware so that a single decision threshold across scores involving various lengths results in poor performance. To remedy this, they propose length-aware score . We use this length-aware score with set to 4; this choice strikes a good balance between generation quality / diversity and robustness to attacks.
Bahri. Bahri and Wieting (2024) recently proposed a distortion-free watermarking scheme that requires only black-box access to an LLM. The scheme works by autoregressively sampling sequences from the LLM, each roughly tokens long. Let be the set of unique sequences from , represent the set of -grams for sequence , and the set of seeds that have had duplicates across the ’s removed and where a random unseen seed is added, if necessary, to ensure all ’s are nonempty. Then is returned, where,
@@
Kirchenbauer. Kirchenbauer et al. (2023) uses the last tokens to pseudorandomly partition the vocabulary for the next token into two lists: a green list of size and a red list consisting of the remainder. A positive bias of is added to the logits of the green list tokens while those of the red list are left unchanged. This has the effect of modifying so that green list tokens are more probable. The score for a text consisting of tokens, of which were found to be green is . We incorporate the latest updates to the algorithm,555https://github.com/jwkirchenbauer/lm-watermarking such as including the current token in the -gram and skipping duplicate -grams at test time. The scheme modifies the model probability so it’s not distortion-free, but a good balance between watermark strength and generation quality can often be achieved. We set , , and .
Kuditipudi. Using the last tokens as the basis for the PRF has the downside that modifying just one of the tokens changes the output and subsequently hurts detection. Kuditipudi et al. (2023) addresses this limitation, which we describe in detail in the Appendix. We follow their methodology and use 256 seeds. To expedite the permutation test, we precompute 5,000 reference values for the secret list. These values are obtained by sampling snippets from the training set of C4-realnewslike (Raffel et al., 2019) across all evaluated target lengths.
@@ 3.4 Detectors
We consider the following competitive 1P and 3P detection strategies. We omit techniques like DetectGPT (Mitchell et al., 2023), Ghostbuster (Verma et al., 2023), and DNA-GPT (Yang et al., 2023b) as the baselines covered here, like RADAR (Hu et al., 2023) and Binoculars (Hans et al., 2024), were shown to outperform them.
Log-Likelihood and Rank. We compute per-token log-likelihood (LLh) and rank features of the target text under a model (taken to be our model, ), as done in prior work (Gehrmann et al., 2019). We note that when scoring a response , standard practice throughout the literature has been to use as the likelihood, where is the null or empty prompt. This is incorrect; the true likelihood is: , where is the distribution over prompts. Since the focus of this work is not on new detection techniques but in novel combinations of existing watermark and detection strategies we continue to define likelihood using the empty prompt. Our features are then:
@@
DetectLLM. Su et al. (2023) proposes a novel combination of zero-shot likelihood and rank features. Specifically, they propose the Log-Likelihood Log-Rank Ratio (LRR) as:
@@
Binoculars. Hans et al. (2024) proposes a 3P detection scheme that leverages uncertainty scores from two LLMs. The Binoculars score is:
@@
RADAR. Hu et al. (2023) proposes an adversarial technique to jointly train an AGC detector with a paraphraser. The paraphraser is trained to generate content that will evade the detector while the detector is trained to discern the paraphraser’s outputs. They show the method outperforms baselines on 4 datasets and 8 LLMs. We use their open-source RoBERTa-large based detector.666https://github.com/IBM/RADAR
RoBERTa Classifier. As done in prior work, we fine-tune a RoBERTa (Liu et al., 2019) model for binary classification. To detect , positive samples are random non-watermarked generations under and negative ones are human responses and non-watermarked generations under . To train the classifier, we use the entire train split of eli5-category (one positive and two negative samples for each prompt), fine-tuning RoBERTa-large with Adam (Kingma and Ba, 2014) for 1 epoch using batch size 32, weight decay 0.01 and a linear learning rate schedule with no warm-up and initial rate 5e-5. To make the detector robust to shorter and truncated texts, the train set is augmented by additionally including a truncated version of each example (to half the number of tokens).
@@ 3.5 Hybrid Detection
We explore a number of designs for hybrid detection. For methods that are trainable or require calibration, we use held-out samples from whichever dataset is not used for testing. That is, when evaluating on databricks-dolly-15k we use eli5-category for calibration (and vice-versa). This ensures that the new detector does not overfit to the test distribution. For all non-cascade models below, scikit-learn (Buitinck et al., 2013) (version 1.5.1) is used with the default hyperparameter settings.
@@
3.6 Adversarial Attacks
In practice, users of our model may act adversarially and attempt to evade detection. Following the methodology of Bahri and Wieting (2024), we study how detectability of our approaches degrades under two attack strategies — random token replacement and paraphrasing. While some may argue it does not make sense to evaluate this setting because, quite simply, the text we are trying to detect is no longer text produced from our model (it is from a paraphrasing LLM for example), we include it nonetheless. We corrupt only the positive test samples (i.e. ’s watermarked generations), leaving the negative test ones untouched. The calibration dataset used for fitting parameters or thresholds is corrupted in the same way as the one used for testing.
@@ 3.6 Adversarial Attacks
Paraphrasing. In this attack, the adversary attempts to evade detection by paraphrasing the watermarked text. We use Gemini-1.5-Pro (Team et al., 2024a) to paraphrase each non-truncated watermarked generation using the prompt: “Paraphrase the following: {RESPONSE}”. We skip examples for which Gemini does not return a response, for instance, for safety reasons.
@@ 3.7 Evaluation Criteria
(1) Partial ROC-AUC. We use partial ROC-AUC (pAUC) up to a max FPR, as utilized in Bahri and Wieting (2024). Since cascades (1S and 2S) have multiple thresholds, we sweep over a grid of thresholds (in 1% increments based on percentiles), record the (FPR, TPR) pairs that result from each threshold setting, and then determine the ROC’s Pareto front, which is subsequently used to calculate pAUC using the trapezoidal rule. All other methods use a single decision threshold and for them we check every test datapoint in generating the (FPR, TPR) pairs, as is the standard practice (e.g. sklearn’s implementation). It is crucial to note that the granularity of the threshold grid search for cascades has a significant impact on the metric (especially so when the region of the ROC we are interested in is slim); performance should only increase the more finely grained the search is, and when every datapoint is checked, it should never underperform either detector since both cascades can represent any decision rule involving a single threshold on watermark or non-watermark scores alike. In other words, any losses for 1S and 2S under this metric should be ignored.
@@
All standard errors reported are obtained by bootstrap resampling each class of the test set separately, to ensure that the resampled test datasets remain class-balanced. standard errors are shown in all tables and figures.

### v2

@@ 1 Introduction
The aim of watermarking is to bias the model so that first-party detection becomes more tractable. Most schemes do not modify the LLM’s training procedure but instead inject the watermark signal at inference time as part of the autoregressive decoding loop (aaronson; bahri2024watermark; kuditipudi2023robust; kirchenbauer2023watermark; dathathri2024scalable). Meanwhile, non-watermark detection has mostly been viewed as a binary classification task, with the goal of discriminating one class from another — typically human-written text from text authored by one or more models. Strategies here include training a binary classifier on labeled text samples or computing uncertainty-based scores, such as likelihood, under a specific LLM (zellers2019defending; solaiman2019release; gehrmann2019gltr; su2023detectllm; mitchell2023detectgpt). The focus of our work is on improving first-party detection by intelligently combining watermark- and non-watermark-based detection approaches.
The entropy of an LLM’s distribution over possible responses conditioned on a specific prompt is a key quantity in watermark detection. For example, bahri2024watermark provides a lower-bound on the detection ROC-AUC as a function of the entropy of the sampled next-token distribution. In all schemes, watermark performance improves with more available entropy. As a concrete example, consider two input prompts to a watermarked LLM: (1) “Where is Mount Whitney?” and (2) “Write me a dreamy haiku about the John Muir Trail in California.” The first prompt is a fact-based inquiry whereas the second is more open-ended in nature, so the entropy will likely be higher for the second prompt than for the first and the generated haiku will be easier to detect via the watermark. The role of entropy for non-watermark detection is less studied. We show how non-watermark detectors, when configured in a hybrid setup, can substantially bolster watermarking in low entropy regimes. For example, training a logistic regression model on watermark scores and scores from a RoBERTa-based AGC classifier can boost watermarking test accuracy from 75% to over 95%, a 20% absolute improvement, on the 20% of test prompts that afford the least amount of entropy. We study how these two approaches to detection can be combined effectively for better predictive performance than either part as well as computational advantages, and we recommend practical algorithms for deployment.
@@ 2.1 Watermarking
Although watermarking, sometimes referred to as linguistic steganography, has a storied past, interest in its application to modern large language models grew substantially after the seminal works of kirchenbauer2023watermark and aaronson. Many successful techniques employ pseudo-random functions (PRFs) and cryptographic hashes and are applied token-by-token during the autoregressive decoding loop. kirchenbauer2023watermark uses a PRF to bias the next-token probability distribution so that certain tokens known only to the watermarker are made more probable. The degree of bias is a hyper-parameter that trades off text distortion and watermark strength. aaronson proposes a clever strategy that selects the token in the next-token distribution that has a high PRF value and a high likelihood in a way that makes the process appear distortion-free. kuditipudi2023robust applies a scheme similar to aaronson but to increase robustness to attacks such as paraphrasing, the pseudo-random numbers are determined by cycling through a secret, predetermined sequence of values rather than by -grams. Meanwhile, SynthID (dathathri2024scalable) samples a large number of tokens from the next-token distribution and pits them head-to-head in a series of tournaments, the winner of which is selected as the next token. lee2023wrote adapts kirchenbauer2023watermark’s scheme for the code-generation task by applying the watermark only at decoding steps that have sufficient entropy.
Black-box schemes that only require a way to sample sequences from the LLM have been devised (yang2023watermarking; giboulot2024watermax; chang-etal-2024-postmark; bahri2024watermark). Most notably, bahri2024watermark uses elements from aaronson to construct a general framework for black-box distortion-free watermarking, which can be used effectively when white-box access is available.
Ways of evading, spoofing, and even stealing watermarks have been extensively studied (zhang2023watermarks; gu2023learnability; jovanovic2024watermark; cheng2025revealing). To address the weakness of many schemes to attacks such as token substitution or paraphrasing, watermarking based on semantics or invariant features has also been proposed (liu2023semantic; hou2023semstamp; ren2023robust; yoo2023robust). fernandez2023three tests various watermarking schemes on classical NLP benchmarks and introduces new statistical tests — for example, they suggest skipping duplicate -grams during testing. liang2024waterpark conducts a comprehensive assessment of watermarking strategies, finding that incorporating a non-watermark RoBERTa-based detector like the way we do can help boost robustness to attacks. The purpose of our work is to generalize this focused insight by proposing and then carefully evaluating various hybrid schemes.
@@ 2.2 AGC Detection
A powerful paradigm to detect AGC is to finetune a pretrained LM on sizable datasets for classification (zellers2019defending; solaiman2019release); notably, hu2023radar and tian2023multiscale do so via adversarial and positive-unlabeled (PU) learning respectively, while tay2020reverse applies it to the task of discriminating different LM generators from one another.
A second camp of approaches focuses on capturing statistical patterns of AGC using little to no training data — such as the intrinsic dimensionality of generated text (tulchinskii2024intrinsic), perplexity / rank / log-rank (gehrmann2019gltr; su2023detectllm), perplexity curvature (mitchell2023detectgpt), and -gram patterns (yang2023dna).
mireshghallah2024smaller shows, interestingly, that small LM generators make for better zero-shot detectors than larger ones. mao2024raidar leverages the observation that LLMs change more words when tasked to rewrite human text than to rewrite AGC. zhang2024llm shows that existing detectors are inadequate for detecting AI-revised human-written text, text that a human wrote with assistance from an LLM. Interestingly, russell-etal-2025-people found that humans who use LLMs for writing tasks can outperform many commercial and open source detectors even in the presence of paraphrasing and humanization.
As usage of large language models is still nascent, it remains to be seen precisely how they are used in the wild, be it with good or nefarious intentions. bahri2021generative conducts a large-scale study to this end by running a GPT-2 detector on half a billion web pages.
Ways to evade detection have been studied. For example, both lu2023large and kumarage2023reliable propose a prompting strategy that encourages LLMs to produce more human-looking text, while krishna2024paraphrasing presents a paraphrasing model called DIPPER that can successfully fool detectors.
@@ 3.1 Models, Datasets, Hyperparameters, and Compute
We consider two distinct settings: when is Gemma-7B-instruct (team2024gemma)111https://huggingface.co/google/gemma-7b-it and is Mistral-7B-instruct (jiang2023mistral)222https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1 as well as the reverse assignment.
We use two test datasets: databricks-dolly-15k333https://huggingface.co/datasets/databricks/databricks-dolly-15k (DatabricksBlog2023DollyV2), an open source dataset of instruction-following examples for brainstorming, classification, closed QA, generation, information extraction, open QA, and summarization. Exactly replicating bahri2024watermark, we use prompts from the brainstorming, generation, open QA (i.e. general QA), and summarization categories, whose human responses are at least 50 tokens long (save one example, which was removed because the prompt was extremely long); this amounts of 5,233 prompts total.
We use the training and test splits of eli5-category444https://huggingface.co/datasets/rexarski/eli5_category (eli5_lfqa). Prompts are formed by concatenating the prompt and selftitle fields. Only examples whose prompt field is non-empty and contains a ? are kept — for a total of 83,089 train and 4,855 test samples.
@@ 3.3 Watermarks
Aaronson. At each step, aaronson computes a pseudorandom number for each token as, , where is the preceding -gram, and denotes concatenation. Token is selected, where . At test time, -grams are extracted from the test query and the detection score is given by,
bahri2024watermark notes that is not length-aware so that a single decision threshold across scores involving various lengths results in poor performance. To remedy this, they propose length-aware score . We use this length-aware score with set to 4; this choice strikes a good balance between generation quality / diversity and robustness to attacks.
Bahri. bahri2024watermark recently proposed a distortion-free watermarking scheme that requires only black-box access to an LLM. The scheme works by autoregressively sampling sequences from the LLM, each roughly tokens long. Let be the set of unique sequences from , represent the set of -grams for sequence , and the set of seeds that have had duplicates across the ’s removed and where a random unseen seed is added, if necessary, to ensure all ’s are nonempty. Then is returned, where,
@@
Kirchenbauer. kirchenbauer2023watermark uses the last tokens to pseudorandomly partition the vocabulary for the next token into two lists: a green list of size and a red list consisting of the remainder. A positive bias of is added to the logits of the green list tokens while those of the red list are left unchanged. This has the effect of modifying so that green list tokens are more probable. The score for a text consisting of tokens, of which were found to be green is . We incorporate the latest updates to the algorithm,555https://github.com/jwkirchenbauer/lm-watermarking such as including the current token in the -gram and skipping duplicate -grams at test time. The scheme modifies the model probability so it’s not distortion-free, but a good balance between watermark strength and generation quality can often be achieved. We set , , and .
Kuditipudi. Using the last tokens as the basis for the PRF has the downside that modifying just one of the tokens changes the output and subsequently hurts detection. kuditipudi2023robust addresses this limitation, which we describe in detail in the Appendix. We follow their methodology and use 256 seeds. To expedite the permutation test, we precompute 5,000 reference values for the secret list. These values are obtained by sampling snippets from the training set of C4-realnewslike (2019t5) across all evaluated target lengths.
@@ 3.4 Detectors
We consider the following competitive 1P and 3P detection strategies. We omit techniques like DetectGPT (mitchell2023detectgpt), Ghostbuster (verma2023ghostbuster), and DNA-GPT (yang2023dna) as the baselines covered here, like RADAR (hu2023radar) and Binoculars (hans2024spotting), were shown to outperform them.
Log-Likelihood and Rank. We compute per-token log-likelihood (LLh) and rank features of the target text under a model (taken to be our model, ), as done in prior work (gehrmann2019gltr). We note that when scoring a response , standard practice throughout the literature has been to use as the likelihood, where is the null or empty prompt. This is incorrect; the true likelihood is: , where is the distribution over prompts. Since the focus of this work is not on new detection techniques but in novel combinations of existing watermark and detection strategies we continue to define likelihood using the empty prompt. Our features are then:
@@
DetectLLM. su2023detectllm proposes a novel combination of zero-shot likelihood and rank features. Specifically, they propose the Log-Likelihood Log-Rank Ratio (LRR) as:
@@
Binoculars. hans2024spotting proposes a 3P detection scheme that leverages uncertainty scores from two LLMs. The Binoculars score is:
@@
RADAR. hu2023radar proposes an adversarial technique to jointly train an AGC detector with a paraphraser. The paraphraser is trained to generate content that will evade the detector while the detector is trained to discern the paraphraser’s outputs. They show the method outperforms baselines on 4 datasets and 8 LLMs. We use their open-source RoBERTa-large based detector.666https://github.com/IBM/RADAR
RoBERTa Classifier. As done in prior work, we fine-tune a RoBERTa (liu2019roberta) model for binary classification. To detect , positive samples are random non-watermarked generations under and negative ones are human responses and non-watermarked generations under . To train the classifier, we use the entire train split of eli5-category (one positive and two negative samples for each prompt), fine-tuning RoBERTa-large with Adam (kingma2014adam) for 1 epoch using batch size 32, weight decay 0.01 and a linear learning rate schedule with no warm-up and initial rate 5e-5. To make the detector robust to shorter and truncated texts, the train set is augmented by additionally including a truncated version of each example (to half the number of tokens).
@@ 3.5 Hybrid Detection
We explore a number of designs for hybrid detection. For methods that are trainable or require calibration, we use held-out samples from whichever dataset is not used for testing. That is, when evaluating on databricks-dolly-15k we use eli5-category for calibration (and vice-versa). This ensures that the new detector does not overfit to the test distribution. For all non-cascade models below, scikit-learn (sklearn_api) (version 1.5.1) is used with the default hyperparameter settings.
@@
3.6 Attacks
In practice, users of our model may act adversarially and attempt to evade detection. Following the methodology of bahri2024watermark, we study how detectability of our approaches degrades under two attack strategies — random token replacement and paraphrasing. While more advanced attacks exist, both these attacks are very easy to deploy and furthermore prior work has shown simple paraphrasing to be highly effective.
While some may argue it does not make sense to evaluate this setting because, quite simply, the text we are trying to detect is no longer text produced from our model per se, we include it nonetheless. We corrupt only the positive test samples (i.e. ’s watermarked generations), leaving the negative test ones untouched. The calibration dataset used for fitting parameters or thresholds is corrupted in the same way as the one used for testing.
@@ 3.6 Attacks
Paraphrasing. In this attack, the adversary attempts to evade detection by paraphrasing the watermarked text. We use Gemini-1.5-Pro (team2024gemini) to paraphrase each non-truncated watermarked generation using the prompt: “Paraphrase the following: {RESPONSE}”. We skip examples for which Gemini does not return a response, for instance, for safety reasons.
@@ 3.7 Evaluation Criteria
(1) Partial ROC-AUC. We use partial ROC-AUC (pAUC) up to a max FPR, as utilized in bahri2024watermark. Since cascades (1S and 2S) have multiple thresholds, we sweep over a grid of thresholds (in 1% increments based on percentiles), record the (FPR, TPR) pairs that result from each threshold setting, and then determine the ROC’s Pareto front, which is subsequently used to calculate pAUC using the trapezoidal rule. All other methods use a single decision threshold and for them we check every test datapoint in generating the (FPR, TPR) pairs, as is the standard practice (e.g. sklearn’s implementation). It is crucial to note that the granularity of the threshold grid search for cascades has a significant impact on the metric (especially so when the region of the ROC we are interested in is slim); performance should only increase the more finely grained the search is, and when every datapoint is checked, it should never underperform either detector since both cascades can represent any decision rule involving a single threshold on watermark or non-watermark scores alike. In other words, any losses for 1S and 2S under this metric should be ignored.
@@
It is important to bear in mind that random text snippets (e.g. from wikipedia) are often used as negative samples in other work, whereas our negatives are model or human answers to the same prompts used to generate positive samples. The consequence is that there are far more semantic and syntactic similarities between the two classes in our work, which makes our classification task intrinsically harder. Concretely, positive and negative samples here will have more overlapping -grams, the basis for many watermarking schemes, and so discrimination via the watermark will be more challenging.
@@
All standard errors reported are obtained by bootstrap resampling each class of the test set separately (500 times), to ensure that the resampled test datasets remain class-balanced. standard errors are shown in all tables and figures. In the Appendix, we estimate the sensitivity of our models’ performance on the calibration data, again via bootstrapping.